OpenStudy (anonymous):

The diagonals of an isosceles trapezoid are labeled a and b. If the length of diagonal a is 3x+9 inches and the length of diagonal b is 2x + 16 inches. What is the length of the diagonals? isosceles trapezoid with diagonals from lower left to upper right and lower right to upper left labeled a and b respectively 12 inches 7 inches 20 inches 30 inches The diagonals of an isosceles trapezoid are labeled a and b. If the length of diagonal a is 3x+9 inches and the length of diagonal b is 2x + 16 inches. What is the length of the diagonals? isosceles trapezoid with diagonals from lower left to upper right and lower right to upper left labeled a and b respectively 12 inches 7 inches 20 inches 30 inches @Mathematics

OpenStudy (aroub):

Remember: The diagonals of an isosceles trapezoid are congruent. so, 3x+9=2x=16 3x-2x=-9+16 x=7

OpenStudy (aroub):

=7 inches

OpenStudy (anonymous):

:)

OpenStudy (aroub):

:D